OpenStudy (anonymous):

Find the Slope of each line: y= -2/5 x -1 and 4x+10y=10.

OpenStudy (anonymous):

use y = mx + b

OpenStudy (anonymous):

m is you slope... look closly at the functions...

OpenStudy (anonymous):

the first one s -2/5 for the second one make sure you ut it in its proper form, solve for y...

OpenStudy (thomas5267):

\[ \begin{align*} y&=&-\frac{2}{5}&&x&+&(-1) \\ y&=&m&&x&+&b \end{align*} \] See the similarities?

OpenStudy (anonymous):

Where do you go from there? Just solve? But how?

OpenStudy (thomas5267):

You do not need to solve. The slope is equal to m.

OpenStudy (anonymous):

So for the second slope.. Would you just do 10y=-4x+10?

OpenStudy (thomas5267):

Yes, but you need to divide the whole equation by 10 to get the slope (i.e. \(\frac{10}{10}y=\frac{-4}{10}x+\frac{10}{10}\) )

OpenStudy (anonymous):

Making your slope -4/10?

OpenStudy (thomas5267):

Yes, as you must have y on one side to work, and don't forget to simplify it.

OpenStudy (anonymous):

-2/5

OpenStudy (thomas5267):

Good!

OpenStudy (anonymous):

So if you compare the two slopes both are -2/5 making them parellel?

OpenStudy (thomas5267):

Yes, but I didn't notice that. Great job!

OpenStudy (anonymous):

Thank you so much!